Vanillin, the compound responsible for the flavour vanilla, has a percentage composition of 63.2% C,

5.26% H, and 31.6% O by mass. Determine the empirical formula of vanillin.


1
Expert's answer
2022-06-06T17:04:07-0400

M(C) = 12 g/mol

M(H) = 1 g/mol

M(O) = 16 g/mol

To find the empirical formula, we divide the percentage of each element by its atomic mass. we divide the result by the smallest of the results. if whole numbers come out, this is the empirical formula. if an integer does not appear, multiply each of them by 2; 3, and so on until an integer is formed

C : 63,2/12 = 5,26 /1,975 =2,66 *3=8

H : 5,26/1 = 5,26 /1,975 =2,66 *3=8

O : 31,6/16 = 1,975 /1,975 =1 *3=3


Answer: C8H8O3